A 62.5-kg person is riding in a car moving at 17.5 m/s (assume this is the positive direction) when the car runs into a bridge abutment. This problem will illustrate why the invention of the airbag dramatically improved the safety of automobiles.

Calculate the horizontal component of the average force, in newtons, on the person if he is stopped by an air bag that compresses an average of 15 cm
Answer #1

using 3rd equation of motion

v^2 = 2 a d

17.5^2 = 2* a * 0.15

a = 1020.83 m/s^2

average force

F = ma = 62.5* 1020.83

F = 63802.1 N

=======

Comment before rate in case any doubt, will reply for sure.. goodluck